# Baker–Akhiezer specialisation of joint eigenfunctions for hyperbolic relativistic Calogero–Moser Hamiltonians

### v1

@@ 3.2 The inductive step
We can now use the $N=2$ case as the basis for proofs by induction. More precisely, we consider $N\geq 3$ and assume that Thm. 3.1 along with Props. 3.63.10 hold true when $N$ is replaced by N1N-1. (In the $N=2$ case, the pertinent results have all been established above.)
@@
$\varphi_{N}(x,y)\equiv e^{\frac{2\pi i}{a}y_{N}\sum_{j=1}^{N}x_{j}}\\ \cdot\int_{\underline{\gamma}}dz\frac{\varphi_{N-1}(z,(y_{1}-y_{N},\ldots,y_{N-1}-y_{N}))\prod_{1\leq j\neq k\leq N-1}\prod_{n=0}^{m-1}2\sinh\left(\frac{\pi}{a}(z_{j}-z_{k}-in)\right)}{\prod_{j=1}^{N}\prod_{k=1}^{N-1}\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}\big(x_{j}-z_{k}+\frac{i}{2}(m-2n-1)\big)\big)},$ (3.28)
@@
$C_{v^{\prime};n_{1},\ldots,n_{N-1}}(x)\\ =(-2ia)^{N-1}\frac{W_{N-1}(x_{1}+i(m-2n_{1}-1)/2,\ldots,x_{N-1}+i(m-2n_{N-1}-1)/2)}{\prod_{k=1}^{N-1}\prod_{n^{\prime}\neq n_{k}}2\sinh\big(\frac{i\pi}{a}(n_{k}-n^{\prime})\big)}\\ \cdot\frac{\varphi_{N-1,\nu^{\prime}}(x_{1}+i(m-2n_{1}-1)/2,\ldots,x_{N-1}+i(m-2n_{N-1}-1)/2)}{\prod_{k=1}^{N-1}\prod_{j\neq k}\prod_{n^{\prime}=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{j}-x_{k}+i(n_{k}-n^{\prime})\big)}$ (3.42)
@@
$\varphi_{N,\rho_{N}(m-1)}(x)=\left(\frac{2a(-i)^{m}}{\prod_{n=1}^{m-1}2\sin(\pi n/a\big)}\right)^{N-1}\\ \cdot\frac{W_{N-1}(x)\varphi_{N-1,\rho_{N-1}(m-1)}(x)}{\prod_{k=1}^{N-1}\prod_{j\neq k}\prod_{n^{\prime}=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{j}-x_{k}+i(m-1-n^{\prime})\big)},$ (3.44)
@@
$\begin{split}\varphi_{N,\sigma_{N-1}(\rho_{N}(m-1))}(x)&=-\left(\frac{2a(-i)^{m}}{\prod_{n=1}^{m-1}2\sin(\pi n/a)}\right)^{N(N-1)/2}\\ &\quad\cdot\frac{1}{\prod_{k=1}^{N}\prod_{j=1}^{\min(k,N-1)-1}\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{k}-x_{j}+in\big)}\\ &\quad\cdot\frac{1}{\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{N-1}-x_{N}+in\big)}\\ &=\varphi_{N,\rho_{N}(m-1)}(\sigma_{N-1}(x)).\end{split}$ (3.46)
@@
where $\Gamma_{0}=\varphi_{N,\rho_{N}(m-1)}$ and $\nu>0$ means that $\nu=\sum_{1\leq j<k\leq N}l_{jk}(e_{j}-e_{k})$ with $l_{jk}\geq 0$ and not all equal to zero. From Proposition 3.7, we get that the functions $\varphi_{N}(\sigma(x),y)$, $\sigma\in S_{N}$, form a linearly independent set; and, since the pre-factor $\prod_{j<k}2\sinh(\pi(y_{j}-y_{k}))$ in (3.30) is antiperiodic under shifts $y_{j}\to y_{j}-i$ and the Macdonald operators $D_{N}^{r}$ are $S_{N}$-invariant, it follows that $\varphi_{N}(x,y)$ and $\varphi_{N}(\sigma_{N-1}(x),\sigma_{N-1}(y))$ satisfy the difference equation
$D_{N}(y)\varphi(x,y)=\varphi(x,y)\sum_{j=1}^{N}e^{\frac{2\pi}{a}x_{j}},\ \ \ D_{N}(y)\equiv D^{1}_{N}(q^{i2y};q^{2},q^{2(1-m)}).$ (3.48)
For the corresponding series $f_{N}$, this translates into the difference equation $L_{N}(y)f(x,y)=f(x,y)\sum_{j=1}^{N}e^{\frac{2\pi}{a}x_{j}}$ given by the difference operator
@@
$\begin{split}\varphi_{N}(a,m;x,y)&=\left(\frac{2a(-i)^{m}}{\prod_{n=1}^{m-1}2\sin(\pi n/a)}\right)^{N(N-1)/2}\\ &\quad\cdot\frac{\psi_{N}(\exp(-i\pi/a),m-1;ix,iy)}{\prod_{1\leq j<k\leq N}\prod_{n=-m+1}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{k}-x_{j}+in)\big)}.\end{split}$ (3.51)
@@ Acknowledgements
Financial support from the Swedish Research Council (Project-id 2024-05649_VR) is gratefully acknowledged.

### v2

@@ 3.2 The inductive step
We can now use the $N=2$ case as the basis for proofs by induction. More precisely, we consider $N\geq 3$ and assume that Thm. 3.1 along with Props. 3.63.7, Lemma 3.8 and Props. 3.93.10 hold true when $N$ is replaced by N1N-1. (In the $N=2$ case, the pertinent results have all been established above.)
@@
$\varphi_{N}(x,y)\equiv e^{\frac{2\pi i}{a}y_{N}\sum_{j=1}^{N}x_{j}}\\ \cdot\int_{\underline{\gamma}}dz\prod_{1\leq j<k\leq N-1}\sinh\left(\frac{\pi}{a}(z_{k}-z_{j})\right)\prod_{n=-m+1}^{m-1}2\sinh\left(\frac{\pi}{a}(z_{k}-z_{j}+in)\right)\\ \cdot\frac{\varphi_{N-1}(z,(y_{1}-y_{N},\ldots,y_{N-1}-y_{N}))}{\prod_{j=1}^{N}\prod_{k=1}^{N-1}\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}\big(x_{j}-z_{k}+\frac{i}{2}(m-2n-1)\big)\big)},$ (3.36)
@@
$C_{v^{\prime};n_{1},\ldots,n_{N-1}}(x)\\ =(-ia)^{N-1}\frac{W_{N-1}(x_{1}+i(m-2n_{1}-1)/2,\ldots,x_{N-1}+i(m-2n_{N-1}-1)/2)}{\prod_{k=1}^{N-1}\prod_{n^{\prime}\neq n_{k}}2\sinh\big(\frac{i\pi}{a}(n_{k}-n^{\prime})\big)}\\ \cdot\frac{\varphi_{N-1,\nu^{\prime}}(x_{1}+i(m-2n_{1}-1)/2,\ldots,x_{N-1}+i(m-2n_{N-1}-1)/2)}{\prod_{k=1}^{N-1}\prod_{j\neq k}\prod_{n^{\prime}=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{j}-x_{k}+i(n_{k}-n^{\prime})\big)}$ (3.50)
@@
$\varphi_{N,\rho_{N}(m-1)}(x)=\left(\frac{a(-i)^{m}}{\prod_{n=1}^{m-1}2\sin(\pi n/a\big)}\right)^{N-1}\\ \cdot\frac{W_{N-1}(x)\varphi_{N-1,\rho_{N-1}(m-1)}(x)}{\prod_{k=1}^{N-1}\prod_{j\neq k}\prod_{n^{\prime}=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{j}-x_{k}+i(m-1-n^{\prime})\big)},$ (3.52)
@@
$\begin{split}\varphi_{N,\sigma_{N-1}(\rho_{N}(m-1))}(x)&=-\left(\frac{a(-i)^{m}}{\prod_{n=1}^{m-1}2\sin(\pi n/a)}\right)^{N(N-1)/2}\\ &\quad\cdot\frac{1}{\prod_{k=1}^{N}\prod_{j=1}^{\min(k,N-1)-1}\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{k}-x_{j}+in\big)}\\ &\quad\cdot\frac{1}{\prod_{n=0}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{N-1}-x_{N}+in\big)}\\ &=-\varphi_{N,\rho_{N}(m-1)}(\sigma_{N-1}(x)).\end{split}$ (3.54)
@@
where $\Gamma_{0}=\varphi_{N,\rho_{N}(m-1)}$ and $\nu>0$ means that $\nu=\sum_{1\leq j<k\leq N}l_{jk}(e_{j}-e_{k})$ with $l_{jk}\geq 0$ and not all equal to zero. From Proposition 3.7, we get that the functions $\varphi_{N}(\sigma(x),y)$, $\sigma\in S_{N}$, form a linearly independent set; and, since the pre-factor $\prod_{j<k}2\sinh(\pi(y_{j}-y_{k}))$ in (3.38) is (anti)periodic under shifts $y_{j}\to y_{j}-i$ and the Macdonald operators $D_{N}^{r}$ are $S_{N}$-invariant, it follows that $\varphi_{N}(x,y)$ and $\varphi_{N}(\sigma_{N-1}(x),\sigma_{N-1}(y))$ satisfy a difference equation
$D_{N}(y)\varphi(x,y)=\varphi(x,y)\cdot\text{const.}\sum_{j=1}^{N}e^{\frac{2\pi}{a}x_{j}},\ \ \ D_{N}(y)\equiv D^{1}_{N}(q^{i2y};q^{2},q^{2(1-m)}).$ (3.56)
For the corresponding series $f_{N}$, this translates into a difference equation $L_{N}(y)f(x,y)=f(x,y)\cdot\text{const.}\sum_{j=1}^{N}e^{\frac{2\pi}{a}x_{j}}$ given by the difference operator
@@
$\begin{split}\varphi_{N}(a,m;x,y)&=\left(\frac{a(-i)^{m}}{\prod_{n=1}^{m-1}2\sin(\pi n/a)}\right)^{N(N-1)/2}\\ &\quad\cdot\frac{\psi_{N}(\exp(-i\pi/a),m-1;ix,iy)}{\prod_{1\leq j<k\leq N}\prod_{n=-m+1}^{m-1}2\sinh\big(\frac{\pi}{a}(x_{k}-x_{j}+in)\big)}.\end{split}$ (3.59)
@@ Acknowledgements
I would like to thank N. Belousov for helpful comments. Financial support from the Swedish Research Council (Project-id 2024-05649_VR) is gratefully acknowledged.